# Multi-Class Stackelberg Games for the Co-Design of Networked Systems

### v2

@@ I Introduction
In this paper, we study the co-design problem based on Stackelberg game. The Stackelberg game framework, originally proposed by Heinrich von Stackelberg in [5], models hierarchical decision-making structures involving a leader and one or more followers. The use of leader-follower schemes has been successfully applied to solve several engineering problems, e.g., formation control [18, 19], crowd dynamics modeling [20], hierarchical scheme for price dynamics analysis under the framework of mean-field-type games [21]. In its original formulation, Stackelberg analyzed the interactions between firms with asymmetric decision-making, where the leader commits to a strategy first, and the followers respond optimally to that strategy. This hierarchical, bi-level decision-making framework is well-suited for addressing the co-design problem, as it mirrors the natural structure of infrastructure design (leader) followed by operational control (follower). Accordingly, we adopt a Stackelberg game formulation to model the inter-dependencies between system design and control, thereby providing a structured methodology for deriving optimal co-design solutions.
@@ II Problem Statement
In this paper, we consider a networked engineering system and aim to analyze two main problems using game-theoretical tools. On one hand, we analyze the design of the networked system led by a leader’s layer. At this design stage, some key parameters for the system are determined. On the other hand, we analyze the control design to operate the system led by a followers’ layer. At this design stage, optimal control policies are defined to operate the system subject to the established design parameters by the leaders. There are two classes of decision-makers, i.e., leaders and followers. The set of $L\in\mathbb{N}$ leaders is given by :=$\mathcal{L}:=\{1,\dots,L\}$, and the set of $M\in\mathbb{N}$ followers is given by :=$\mathcal{M}:=\{1,\dots,M\}$. The leader decision-makers set takes care of the system design and parameter setting, whereas the second decision-makers set is in charge of designing the optimal control to operate the system, adhering to the imposed rules of the leader’s set. The leaders interact with each other using a static normal form game, and their decisions are coupled with the followers’ interactions. The followers interact with each other by means of a difference game whose settings are leader-strategic-dependent.
@@ II-A Followers Strategic Interaction and Settings
The followers interact on a difference game problem within a discrete-time interval $[0..T]:=[0,T]\cap\mathbb{Z}_{+}$, with $T\in\mathbb{N}$, involving a dynamical system given as follows:
@@
[…]
@@
$\mathbf{u}_{-i}$ $:=(\mathbf{u}_{1},\dots,\mathbf{u}_{i-1},\mathbf{u}_{i+1},\dots,\mathbf{u}_{M})\in\mathcal{U}=\prod_{j\in\mathcal{M}\setminus\{i\}}\mathcal{U}_{j},$ (2)
@@ II-C Followers Cooperative Behavior
s. t. {(1),$\ x_{k}\in\mathbb{X}(\{\text{leader strategies}\}),\\ \text{for all}~{}k\in[0..T],\end{cases}$ (6)
@@ II-F Leaders Cooperative Behavior
For notation simplicity, let $J(\mathbf{a},\{\text{follower strategies}\})$ be as
@@ II-G Price of Anarchy for Leaders and Followers
Note that the denominators in both expressions for the price of anarchy are computed by solving the cooperative game problem presented in Problem (10) and Problem (5). Therefore, the denominators become $J(\tilde{\mathbf{a}}^{*},\{\text{follower strategies}\})$ and $V(x_{0},\tilde{\mathbf{u}}^{*})$, respectively. In addition, let us assume that, as we will see in the numerical example we present in this paper, there is a unique Nash equilibrium for each of the games across the layers, i.e. ${\mathbf{a}}^{*}$ and ${\mathbf{u}}^{*}$. Then, the prices of anarchy can be written as
@@
Class II $(\tilde{\mathbf{a}}^{*},\tilde{\mathbf{u}}^{*})$ Cooperative Cooperative
Class III $(\mathbf{a}^{*},\tilde{\mathbf{u}}^{*})$ Non-cooperative Cooperative
@@ III-A Definition of Stackelberg Equilibrium Classes
$a_{j}^{*}\in\arg\min_{a_{j}\in A_{j}}J_{j}\left(a_{j},a_{-j}^{*},\mathbf{u}^{*}\bigg{|}\begin{array}[]{l}\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},a_{j},a_{-j}^{*})\\ \forall i\in\mathcal{M}\end{array}\right).$ (12e)
@@
$\tilde{\mathbf{u}}^{*}$ $\in\mathbb{BR}^{\mathcal{M}}(\tilde{\mathbf{a}}^{*}),$ (13a)
$\tilde{\mathbf{a}}^{*}$ $\in\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{M}}J_{j}(\mathbf{a},\tilde{\mathbf{u}}^{*}|\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(\tilde{\mathbf{a}})).$ (13b)
@@
$\forall~{}i\in\mathcal{M}:~{}\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},\tilde{\mathbf{a}}^{*}),$
@@
$\forall~{}j\in\mathcal{L}:a_{j}^{*}\in\arg\min_{a_{j}\in A_{j}}J_{j}(a_{j},a_{-j}^{*},\tilde{\mathbf{u}}^{*}|\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(a_{j},a_{-j}^{*})).$
@@ III-B1 Class I: Stackelberg game with non-cooperative leaders and non-cooperative followers
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$ $A_{j},~{}\forall j\in\mathcal{L}$,
@@ III-B2 Class II: Stackelberg game with cooperative leaders and cooperative followers
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$, $A_{j},~{}\forall j\in\mathcal{L}$
@@ III-B3 Class III: Stackelberg game with non-cooperative leaders and cooperative followers
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$, $A_{j},~{}\forall j\in\mathcal{L}$
@@ III-B4 Class IV: Stackelberg game with cooperative leaders and non-cooperative followers
The computation of the Stackelberg equilibrium for this setting proceeds as follows. First, for every feasible leader strategic profile $\mathbf{a}\in\mathcal{A}$, the followers solve a non-cooperative dynamic game to compute the corresponding Nash equilibrium $\mathbf{u}^{*}(\mathbf{a})\in\mathbb{BR}^{\mathrm{NC}}_{\mathcal{M}}(\mathbf{a})$, which represents the followers’ best responses to the leaders’ decisions. Second, using the equilibrium strategic profile $\mathbf{u}^{*}(\mathbf{a})$ in terms of the leaders’ selections, the leader cooperative problem can be stated only in terms of $\mathbf{a}$, and the joint optimization for the leaders can be computed, i.e., $\tilde{\mathbf{a}}^{*}\in\mathcal{A}$ can be found. Then, using the cooperative solution for the leaders $\tilde{\mathbf{a}}^{*}$, the Nash equilibrium for the follower game problem becomes $\mathbf{u}^{*}:=\mathbf{u}^{*}(\tilde{\mathbf{a}}^{*})$. The resulting pair $(\mathbf{u}^{*},\tilde{\mathbf{a}}^{*})$ constitutes the Class IV Stackelberg equilibrium, integrating the leaders’ cooperative optimization with the followers’ strategic competition. The complete procedure for this computation is detailed in Algorithm III-B4.
@@
=$\mathcal{L}=\{1,\dots,L\}$, =$\mathcal{M}=\{1,\dots,M\}$, $A_{j},~{}\forall j\in\mathcal{L}$
@@ IV Networked System Application
The Barcelona Drinking Water Network (DWN) in [30, Chapter 3, Fig. 3.4], managed by the company Aguas de Barcelona (AGBAR), supplies drinking water to Barcelona city and its metropolitan area, utilizing water from the Ter and Llobregat rivers—regulated upstream by dams with a combined reservoir capacity of 600 hm³, alongside groundwater from the Besós River aquifer and supplementary wells. The network integrates four drinking water treatment plants: the Abrera and Sant Joan Despì facilities (Llobregat River), the Cardedeu plant (Ter River), and the Besós plant (groundwater), with additional pumping infrastructure extracting from wells to achieve a total flow of approximately 7 m³/s. This is a well-known benchmark that has been used to illustrate networked control applications, game-theory-based control analysis, and optimization-based controllers. Here, we use this case study to evaluate and illustrate the Stackelberg game classes introduced in this paper.
@@
${x}_{k+1}$ $=A{x}_{k}+\sum_{i\in\mathcal{M}}B_{i}{u}_{i,k}+B_{l}{d}_{k},~{}\forall k\in[0..(T-1)],$ (14a)
@@
$\mathbb{X}(\cdot)$ =$=\{{x}\in\mathbb{R}^{17}:{x}^{\mathrm{min}}\leq{x}\leq{x}^{\mathrm{max}}(\cdot)\},$
@@ IV-A Leaders and Followers in the Barcelona DWN
In terms of the co-design problem of the Barcelona DWN, we consider $L=4$ leaders within the system in charge of the design of four reservoirs’ dimensions, i.e., =$\mathcal{L}=\{1,\dots,4\}$ corresponding to the states $x_{4},x_{10},x_{14},x_{1}$, respectively (see Table II). Each leader in $\mathcal{L}$ has a set of possible actions in the strategic game
@@ 4 Red 2 $x_{1}$
On the other hand, let us consider $M=4$ followers, i.e., =$\mathcal{M}=\{1,\dots,4\}$, which are in charge of four different sub-systems of the DWN. As shown in Fig. 2, four subsystems highlighted by different colors are considered. This partitioning has been adopted from the research reported in [30, Chapter 8, Fig. 8.3]. Thus, each follower decides the control inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{4}$ corresponding to the sub-systems with colors green, blue, magenta, and red, respectively.
@@ IV-C Non-Cooperative and Cooperative Followers Game
[…]
@@
$\mathbb{U}_{i}$ :=$:=\{{u}_{i}\in\mathbb{R}^{n_{u_{i}}}:{u}^{\mathrm{min}}_{i}\leq{u}_{i}\leq{u}^{\mathrm{max}}_{i}\},$ (18)
$\mathbb{X}(\mathbf{a})$ :=$:=\{{x}\in\mathbb{R}^{n_{x}}:{x}^{\mathrm{min}}\leq{x}\leq{x}^{\mathrm{max}}(\mathbf{a})\}.$ (19)
@@
[…]
@@ V Results and Discussions
Leader Follower $\mathbf{a}^{*}$ $J_{1}^{*}$ $J_{2}^{*}$ $J_{3}^{*}$ $J_{4}^{*}$ $J^{*}$
Class I Non-cooperative Non-cooperative [0.5, 0.5, 0.5, 1]0.50.50.51[0.5,\;0.5,\;0.5,\;1][ 0.5 , 0.5 , 0.5 , 1 ] 252,647 252,711 252,819 252,635 1,010,812
Class II Cooperative Cooperative [1.5, 1.25, 0.5, 1.5]1.51.250.51.5[1.5,\;1.25,\;0.5,\;1.5][ 1.5 , 1.25 , 0.5 , 1.5 ] 250,340 250,491 250,479 250,297 1,001,607
Class III Non-cooperative Cooperative [0.5, 0.5, 0.5, 0.5]$0.5$[0.5,\;0.5,\;0.5,\;0.5][ 0.5 , 0.5 , 0.5 , 0.5 ] 250,386 250,449 250,558 250,372 1,001,765
Class IV Cooperative Non-cooperative [1, 0.75, 0.5, 1]10.750.51[1,\;0.75,\;0.5,\;1][ 1 , 0.75 , 0.5 , 1 ] 252,620 252,707 252,776 252,592 1,010,695
@@
Leader Follower $\mathbf{a}^{*}$ $V_{1}^{*}$ $V_{2}^{*}$ $V_{3}^{*}$ $V_{4}^{*}$ $V^{*}$
Class I Non-cooperative Non-cooperative [0.5, 0.5, 0.5, 1]0.50.50.51[0.5,\;0.5,\;0.5,\;1][ 0.5 , 0.5 , 0.5 , 1 ] 0.0732 17.4186 18.3429 0.7798 36.6145
Class II Cooperative Cooperative [1.5, 1.25, 0.5, 1.5]1.51.250.51.5[1.5,\;1.25,\;0.5,\;1.5][ 1.5 , 1.25 , 0.5 , 1.5 ] 0.0743 17.0877 17.5127 1.7343 36.4090
Class III Non-cooperative Cooperative [0.5, 0.5, 0.5, 0.5]$0.5$[0.5,\;0.5,\;0.5,\;0.5][ 0.5 , 0.5 , 0.5 , 0.5 ] 0.0740 17.0852 17.5375 1.7099 36.4066
Class IV Cooperative Non-cooperative [1, 0.75, 0.5, 1]10.750.51[1,\;0.75,\;0.5,\;1][ 1 , 0.75 , 0.5 , 1 ] 0.0716 17.3884 18.3939 0.7383 36.5923
@@
[…]

### v3

@@ I Introduction
In this paper, we study the co-design problem based on Stackelberg game. The Stackelberg game framework, originally proposed by Heinrich von Stackelberg in [5], models hierarchical decision-making structures involving a leader and one or more followers. The use of leader-follower schemes has been successfully applied to solve several engineering problems, e.g., formation control [18, 19], crowd dynamics modeling [20], hierarchical scheme for price dynamics analysis under the framework of mean-field-type games [21], consensus for nonlinear multi-agent systems [22]. In its original formulation, Stackelberg analyzed the interactions between firms with asymmetric decision-making, where the leader commits to a strategy first, and the followers respond optimally to that strategy. This hierarchical, bi-level decision-making framework is well-suited for addressing the co-design problem, as it mirrors the natural structure of infrastructure design (leader) followed by operational control (follower). Accordingly, we adopt a Stackelberg game formulation to model the inter-dependencies between system design and control, thereby providing a structured methodology for deriving optimal co-design solutions.
@@ II Problem Statement
In this paper, we consider a networked engineering system and aim to analyze two main problems using game-theoretical tools. On one hand, we analyze the design of the networked system led by a leader’s layer. At this design stage, some key parameters for the system are determined. On the other hand, we analyze the control design to operate the system led by a followers’ layer. At this design stage, optimal control policies are defined to operate the system subject to the established design parameters by the leaders. There are two classes of decision-makers, i.e., leaders and followers. The set of $L\in\mathbb{N}$ leaders is given by $\mathcal{L}:=\{1,\dots,L\}$, and the set of $M\in\mathbb{N}$ followers is given by $\mathcal{M}:=\{1,\dots,M\}$. The leader decision-makers set takes care of the system design and parameter setting, whereas the second decision-makers set is in charge of designing the optimal control to operate the system, adhering to the imposed rules of the leader’s set. The leaders interact with each other using a static normal form game, and their decisions are coupled with the followers’ interactions. The followers interact with each other by means of a difference game whose settings are leader-strategic-dependent.
@@ II-A Followers Strategic Interaction and Settings
The followers interact on a difference game problem within a discrete-time interval $0..T]:=[0,T]\cap\mathbb{Z}_{+}$, with $T\in\mathbb{N}$, involving a dynamical system given as follows:
@@
where $x\in\mathbb{X}(\cdot)\subset\mathbb{R}^{n_{x}}$ denotes the system states. We highlight that the feasible set for the system states depends on a given parameters, which comes from the leader’s strategic selection, i.e., $\mathbb{X}(\{\text{leader strategies}\})$. The strategic selection of the $i$-th follower decision-maker is denoted by $u_{i}\in\mathbb{U}_{i}\subset\mathbb{R}^{n_{u_{i}}}$. Let
@@
$\mathbf{u}_{-i}$ $:=(\mathbf{u}_{1},\dots,\mathbf{u}_{i-1},\mathbf{u}_{i+1},\dots,\mathbf{u}_{M})\in\mathcal{U}=\hskip-5.69046pt\prod_{j\in\mathcal{M}\setminus\{i\}}\mathcal{U}_{j},$ (2)
@@ II-C Followers Cooperative Behavior
[…]
@@ II-F Leaders Cooperative Behavior
[…]
@@ II-G Price of Anarchy for Leaders and Followers
[…]
@@
[…]
@@ III-A Definition of Stackelberg Equilibrium Classes
$a_{j}^{*}\in\arg\min_{a_{j}\in A_{j}}J_{j}\left(a_{j},a_{-j}^{*},\mathbf{u}^{*}\bigg{|}\begin{array}[]{l}\mathbf{u}_{i}^{*}\in\mathbb{BR}^{\mathcal{M}}_{i}(\mathbf{u}_{-i}^{*},a_{j},a_{-j}^{*})\\ \forall i\in\mathcal{M}\end{array}\hskip-5.69046pt\right).$ (12e)
@@
[…]
$\tilde{\mathbf{a}}^{*}$ $).\displaystyle\in\arg\min_{\mathbf{a}\in\mathcal{A}}\sum_{j\in\mathcal{M}}J_{j}(\mathbf{a},\tilde{\mathbf{u}}^{*}|\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(\tilde{\mathbf{a}}$)). (13b)
@@
[…]
@@
$\forall~j\in\mathcal{L}:a_{j}^{*}\in\arg\min_{a_{j}\in A_{j}}\hskip-2.84544ptJ_{j}(a_{j},a_{-j}^{*},\tilde{\mathbf{u}}^{*}|\tilde{\mathbf{u}}^{*}\in\mathbb{BR}^{\mathcal{M}}(a_{j},a_{-j}^{*})).$
@@ III-B1 Class I: Stackelberg game with non-cooperative leaders and non-cooperative followers
$\mathcal{L}=\{1,\dots,L\}$, $\mathcal{M}=\{1,\dots,M\}$ $A_{j},~\forall j\in\mathcal{L}$,
@@ III-B2 Class II: Stackelberg game with cooperative leaders and cooperative followers
$\mathcal{L}=\{1,\dots,L\}$, $\mathcal{M}=\{1,\dots,M\}$, $A_{j},~\forall j\in\mathcal{L}$
@@ III-B3 Class III: Stackelberg game with non-cooperative leaders and cooperative followers
$\mathcal{L}=\{1,\dots,L\}$, $\mathcal{M}=\{1,\dots,M\}$, $A_{j},~\forall j\in\mathcal{L}$
@@ III-B4 Class IV: Stackelberg game with cooperative leaders and non-cooperative followers
[…]
@@
$\mathcal{L}=\{1,\dots,L\}$, $\mathcal{M}=\{1,\dots,M\}$, $A_{j},~\forall j\in\mathcal{L}$
@@ IV Networked System Application
The Barcelona Drinking Water Network (DWN) in [31, Chapter 3, Fig. 3.4], managed by the company Aguas de Barcelona (AGBAR), supplies drinking water to Barcelona city and its metropolitan area, utilizing water from the Ter and Llobregat rivers—regulated upstream by dams with a combined reservoir capacity of 600 hm³, alongside groundwater from the Besós River aquifer and supplementary wells. The network integrates four drinking water treatment plants: the Abrera and Sant Joan Despì facilities (Llobregat River), the Cardedeu plant (Ter River), and the Besós plant (groundwater), with additional pumping infrastructure extracting from wells to achieve a total flow of approximately 7 m³/s. This is a well-known benchmark that has been used to illustrate networked control applications, game-theory-based control analysis, and optimization-based controllers. Here, we use this case study to evaluate and illustrate the Stackelberg game classes introduced in this paper.
@@
${x}_{k+1}$ $=A{x}_{k}+\hskip-2.84544pt\sum_{i\in\mathcal{M}}\hskip-2.84544ptB_{i}{u}_{i,k}+B_{l}{d}_{k},~\forall k\in[0..(T-1)],$ (14a)
@@
$\mathbb{X}(\cdot)$ $=\{{x}\in\mathbb{R}^{17}:{x}^{\mathrm{min}}\leq{x}\leq{x}^{\mathrm{max}}(\cdot)\},$
@@ IV-A Leaders and Followers in the Barcelona DWN
In terms of the co-design problem of the Barcelona DWN, we consider $L=4$ leaders within the system in charge of the design of four reservoirs’ dimensions, i.e., $\mathcal{L}=\{1,\dots,4\}$ corresponding to the states $x_{4},x_{10},x_{14},x_{1}$, respectively (see Table II). Each leader in $\mathcal{L}$ has a set of possible actions in the strategic game
@@ 4 Red 2 $x_{1}$
On the other hand, let us consider $M=4$ followers, i.e., $\mathcal{M}=\{1,\dots,4\}$, which are in charge of four different sub-systems of the DWN. As shown in Fig. 2, four subsystems highlighted by different colors are considered. This partitioning has been adopted from the research reported in [31, Chapter 8, Fig. 8.3]. Thus, each follower decides the control inputs $\mathbf{u}_{1},\dots,\mathbf{u}_{4}$ corresponding to the sub-systems with colors green, blue, magenta, and red, respectively.
@@ IV-C Non-Cooperative and Cooperative Followers Game
$\text{s.~t.}\begin{cases}{x}_{k+1}=A{x}_{k}+\sum_{i\in\mathcal{M}}B_{i}{u}_{i,k}+B_{l}{d}_{k},\\ 0=\sum\limits_{i\in\mathcal{M}}E_{i}{u}_{i,k}+E_{d}{d}_{k},\\ {u}_{i,k}\in\mathbb{U}_{i},~i\in\mathcal{M},\\ {x}_{k}\in\mathbb{X}(\mathbf{a}),\end{cases}$
@@
$\mathbb{U}_{i}$ $:=\{{u}_{i}\in\mathbb{R}^{n_{u_{i}}}:{u}^{\mathrm{min}}_{i}\leq{u}_{i}\leq{u}^{\mathrm{max}}_{i}\},$ (18)
$\mathbb{X}(\mathbf{a})$ $:=\{{x}\in\mathbb{R}^{n_{x}}:{x}^{\mathrm{min}}\leq{x}\leq{x}^{\mathrm{max}}(\mathbf{a})\}.$ (19)
@@
$\text{s.~t.}~\begin{cases}{x}_{k+1}=A{x}_{k}+B{u}_{k}+B_{l}{d}_{k},\\ 0=E{u}_{k}+E_{d}{d}_{k},\\ {u}_{i,k}\in\mathbb{U}_{i},~i\in\mathcal{M},\\ {x}_{k}\in\mathbb{X}(\mathbf{a}),\end{cases}$
@@ V Results and Discussions
Leader Follower $\mathbf{a}^{*}$ $J_{1}^{*}$ $J_{2}^{*}$ $J_{3}^{*}$ $_{4}^{*} JJ^{*}$
Class I Non-cooperative Non-cooperative [0.5, 0.5, 0.5, 1][0.5,\;0.5,\;0.5,\;1] 252,647 252,711 252,819 252,635 1,010,812
Class II Cooperative Cooperative [1.5, 1.25, 0.5, 1.5][1.5,\;1.25,\;0.5,\;1.5] 250,340 250,491 250,479 250,297 1,001,607
Class III Non-cooperative Cooperative [0.5, 0.5, 0.5, 0.5][0.5,\;0.5,\;0.5,\;0.5] 250,386 250,449 250,558 250,372 1,001,765
Class IV Cooperative Non-cooperative [1, 0.75, 0.5, 1][1,\;0.75,\;0.5,\;1] 252,620 252,707 252,776 252,592 1,010,695
@@
Leader Follower $\mathbf{a}^{*}$ $V_{1}^{*}$ $V_{2}^{*}$ $V_{3}^{*}$ $_{4}^{*} VV^{*}$
Class I Non-cooperative Non-cooperative [0.5, 0.5, 0.5, 1][0.5,\;0.5,\;0.5,\;1] 0.0732 17.4186 18.3429 0.7798 36.6145
Class II Cooperative Cooperative [1.5, 1.25, 0.5, 1.5][1.5,\;1.25,\;0.5,\;1.5] 0.0743 17.0877 17.5127 1.7343 36.4090
Class III Non-cooperative Cooperative [0.5, 0.5, 0.5, 0.5][0.5,\;0.5,\;0.5,\;0.5] 0.0740 17.0852 17.5375 1.7099 36.4066
Class IV Cooperative Non-cooperative [1, 0.75, 0.5, 1][1,\;0.75,\;0.5,\;1] 0.0716 17.3884 18.3939 0.7383 36.5923
@@
Regarding the strategic selection for the leaders, this is presented in Tables III and IV. The Nash equilibrium for the leaders correspond to $\mathbf{a}^{*}=[0.5,\;0.5,\;0.5,\;1]$ and $\mathbf{a}^{*}=[0.5,\;0.5,\;0.5,\;0.5]$ for the Class I and Class III, respectively. Interestingly, we see many similarities between both strategic profiles. Note that, when changing the strategic behavior of the followers from non-cooperative to cooperative, only one of the leaders deviates from its strategic selection. For the cooperative scenario for the leaders, we observe that the optimal solutions are $\mathbf{a}*=[1.5,\;1.25,\;0.5,\;1.5]$ and $\mathbf{a}*=[1,\;0.75,\;0.5,\;1]$ exhibiting a strategic deviation for all the leaders, except one, when the followers change their behavior from non-cooperative to cooperative.